# Beyond Pairwise: Empowering LLM Alignment With Ranked Choice Modeling

### v1

@@ Abstract
Alignment of large language models (LLMs) has predominantly relied on pairwise preference optimization, where annotators select the better of two responses to a prompt. While simple, this approach overlooks the opportunity to learn from richer forms of human feedback, such as multiwise comparisons and top-$k$ rankings. We propose Ranked Choice Preference Optimization (RCPO), a unified framework that bridges preference optimization with (ranked) choice modeling via maximum likelihood estimation. The framework is flexible, supporting both utility-based and rank-based choice models. It subsumes several existing pairwise methods (e.g., DPO, SimPO), while providing principled training objectives for richer feedback formats. We instantiate this framework with two representative ranked choice models (Multinomial Logit and Mallows-RMJ). Empirical studies on Llama-3-8B-Instruct and Gemma-2-9B-it across AlpacaEval 2 and Arena-Hard benchmarks show that RCPO consistently outperforms competitive baselines. RCPO shows how directly leveraging ranked preference data, combined with the right choice models, yields more effective alignment. It offers a versatile and extensible foundation for incorporating (ranked) choice modeling into LLM training.
@@ 1 Introduction
Large language models (LLMs), a prominent form of generative AI, have rapidly transformed human-computer interaction, powering applications from open-ended dialogue (Thoppilan et al., 2022) and content creation (Brown et al., 2020) to code generation (Chen et al., 2021; Li et al., 2023) and healthcare decision support (Nori et al., 2023). A key driver of this success is alignment—training models to produce outputs that are factual, helpful, safe, and aligned with social norms.
@@
Despite their differences, most alignment methods rely on pairwise preference data, where for each prompt $x$, a preferred response $y_{w}$ and a dispreferred response $y_{l}$ are selected by human annotators or AI judges. In practice, however, preference feedback is often richer than simple pairs. Annotators may provide partial rankings, top-$k$ selections, or single-best judgments from a larger candidate set. Current approaches typically reduce this richer information to pairs—e.g., in training InstructGPT, Ouyang et al. (2022) collected rankings of $K$ responses per prompt but converted them into all $\binom{K}{2}$ pairs; in many academic alignment studies (Meng et al., 2024; Chen et al., 2025; Zhao et al., 2024; Gupta et al., 2025), multiple responses are scored by a reward model, but only the highest- and lowest-scoring are kept. These reductions, while convenient for pairwise-based algorithms, can distort the original preference structure and discard potentially valuable information.
@@
To address this gap, we propose Ranked Choice Preference Optimization (RCPO), a general framework that generalizes pairwise preference alignment to ranked choice feedback (see Figure 1). Rather than restricting evaluators to comparing only two responses, RCPO presents a set of candidates and asks them to choose either the single-best or the top-$k$ responses for a given prompt. Our approach is grounded in the theory of (ranked) choice models, in particular discrete choice modeling, which have been extensively studied in psychology, marketing, economics, and operations research. To the best of our knowledge, RCPO is among the first systematic applications of (ranked) choice modeling to LLM alignment.
@@
(1) Conceptual Framework: We establish a systematic connection between LLM fine-tuning and choice modeling, showing that fine-tuning can be essentially reduced to maximum likelihood estimation (MLE) of choice models. Building on this insight, we develop RCPO as a principled extension of pairwise preference alignment that directly incorporates ranked choice feedback. RCPO is a general and flexible framework: any choice model that satisfies certain regularity conditions can be integrated. By preserving the richness of the original annotations – whether single-best or top-$k$ preferences – RCPO avoids the information loss inherent in pairwise conversion and enables more faithful alignment with human intent.
(2) Concrete Examples: We showcase how two broad classes of choice models, i.e., utility- or rank-based, can be accommodated within the RCPO framework. We then instantiate RCPO with a representative model from each class: the Multinomial Logit (MNL) model (McFadden, 1972) for utility-based choices and the Mallows-RMJ model (Feng & Tang, 2022) for rank-based choices. For both models, we derive alignment objectives under single-best and top-$k$ settings (see Table 1). We analyze the gradients to provide theoretical insight into how these objectives drive preference optimization.
@@
(3) Experiments: We evaluate RCPO on state-of-the-art instruction-tuned LLMs (Llama-3-8B-instruct and Gemma-2-9B-it) using widely adopted benchmarks (AlpacaEval 2 and Arena-Hard). Results consistently show that RCPO improves model performance and demonstrates flexibility across base models, preference feedback, and evaluation settings. We highlight the Mallows-RMJ-based preference optimization, which achieves strong results under both pairwise and ranked-choice setups.
@@
2 A Choice-Based Alignment Framework
@@ 2.1 RLHF and DPO
RLHF comprises three sequential phases. First, it fine-tunes a pre-trained LLM through supervised learning on curated instruction data to obtain a reference model $\pi_{\mathrm{ref}}$. Second, RLHF fits a reward model $r^{*}(x,y)$, which can be a neural network itself, based on a separate pairwise preference dataset $\mathcal{D}=\{(x^{(i)},\,y_{w}^{(i)},\,y_{l}^{(i)})\}_{i=1}^{N}$. Here $x^{(i)}$ represents the prompt provided to annotator $i$, and $y_{w}^{(i)}$ and $y_{l}^{(i)}$ are the preferred and dispreferred responses, respectively. Third, the LLM is further fine-tuned via reinforcement learning to maximize the regularized expected reward:
@@
Discrete Choice Models. We also introduce the key concepts for (ranked) choice modeling. First, a rich body of research across economics, marketing, and operations research has developed a variety of choice models to represent human preferences. Let $\mathcal{N}=\{y_{1},y_{2},\ldots,y_{n}\}$ denote the universe of items. A discrete choice model specifies the probability of selecting an item $y$ from an assortment $S\subseteq\mathcal{N}$ under a given context $x$, denoted by $\mathbb{P}(y\mid S\,;x)$. This is an abstraction of many business and economics use cases. For instance, in retail settings, items typically correspond to products, assortments represent the sets of available products at the time of choice, and the context encompasses covariates such as prices, promotions, or product features. There are many ways to define a discrete choice model, which essentially reduce to specifying the probability function $\mathbb{P}(y\mid S\,;x)$.
Ranked Choice Models. Discrete choice models can be extended to richer feedback in the form of ranked choices, which is first formally studied by Feng & Tang (2022) to the best of our knowledge. Let $\mu^{k}=y_{1}\succ y_{2}\succ\cdots\succ y_{k}$ with $\{y_{1},\ldots,y_{k}\}\subseteq S$ denote a top-$k$ list of items from $S$. A ranked choice model defines $\mathbb{P}(\mu^{k}\mid S\ ;x)$, which specifies the probability of observing the partial ranking $\mu^{k}$ under context $x$. This notion of ranked choices generalizes several common feedback structures, such as (i) pairwise comparisons; (ii) discrete choices (or multi-wise comparisons); (iii) listwise feedback (or full rankings over the items in any given assortment); (iv) top-$k$ rankings over a full item set, to name a few. For example, when $k=1$, a ranked choice model reduces to a discrete choice model.
@@ 2.3 RCPO: Connection between LLM Alignment and Ranked Choice Modeling
What truly demonstrates the versatility of RCPO is how it gives birth to new preference optimization methods. In this paper, we consider two prominent classes of choice models: utility-based models, which rely on the numerical magnitudes of the items’ utilities, and rank-based models, which depend on the rankings of the items.
@@ 2.4 Utility-based Choice Models
RUMs are categorized by the distribution of their stochastic terms. The multinomial logit (MNL) model, introduced by McFadden (1972), assumes i.i.d. Gumbel noise. Alternatives include the probit model (joint normal distribution (Daganzo, 2014)), the nested logit model (correlated extreme value distributions (McFadden, 1980)), and the exponomial model (negative exponential distributions (Alptekinoğlu & Semple, 2016)).
@@ 2.5 Rank-based Choice Models
While utility-based models operate on numerical utilities, rank-based models represent preferences as complete orderings over $\mathcal{N}$, depending only on the relative positions of items. See Jagabathula & Venkataraman (2022) for a survey of such models.
The Mallows-type model (Mallows, 1957; Fligner & Verducci, 1986) is among the most widely used classes of ranking models. Let $\mathfrak{S}_{n}$ be the set of permutations of $\mathcal{N}$. A Mallows-type model assigns a probability distribution over permutations $\mu\in\mathfrak{S}_{n}$ based on their distance from a central ranking $\mu_{0}$:
@@
where $d(\cdot,\cdot)$ is a distance function between permutations, $\phi\in(0,1)$ is a dispersion parameter. Intuitively, rankings closer to $\mu_{0}$ are exponentially more likely. Different distance choices yield different variants, such as Kendall’s Tau (Mallows, 1957), Spearman’s rank and footrule (Diaconis & Graham, 1977), Hamming distance (Bookstein et al., 2002), Cayley distance (Irurozki et al., 2018), and Reverse Major Index (Feng & Tang, 2022). To capture context dependence, we parameterize the central ranking and dispersion as functions of $x$, denoted $\mu_{0}(\cdot|x)$ and $\phi(x)$. For readability, we may suppress $x$ in the notation when its dependence is unambiguous.
@@ 3.1 Multinomial Logit Model (MNL)
If we take the random shock $\boldsymbol{\varepsilon}$ to be i.i.d. Gumbel, we recover the Multinomial Logit (MNL) model (McFadden, 1972), arguably the most widely used RUM. The corresponding choice probabilities in 4 admit simple closed-form expressions, given by $\mathbb{P}(y\mid S\ ;x)=e^{\nu_{y}(x)}/\sum_{y^{\prime}\in S}e^{\nu_{y^{\prime}}(x)}$. As such, it naturally extends the Bradley-Terry model from pairwise comparisons to multi-item choice settings. By representing the mean utility as the reward defined in 2, we have the following theorem.
@@
The MNL model also extends to a simple ranked choice model (Luce et al., 1959; Plackett, 1975; Feng & Tang, 2023). Specifically, the probability that a top-$k$ ranked choice $\mu^{k}=y_{1}\succ y_{2}\succ\ldots\succ y_{k}$ is chosen out of an assortment $S\subseteq\mathcal{N}$ is given by
$\mathbb{P}(\mu^{k}\mid S\ ;x)=\prod\nolimits_{i=1}^{k}\tfrac{e^{\nu_{y_{i}}(x)}}{\sum_{y\in S\setminus\{y_{1},\ldots,y_{i-1}\}}e^{\nu_{y}(x)}}.$ (8)
@@ 3.2 Mallows-RMJ Model
A notable challenge in applying Mallows-type models to (ranked) choice modeling is that the ranked choice probabilities in (6) are usually difficult to obtain, since the sum is taken over all permutations. In this regard, an exception is Feng & Tang (2022), who adopt a Mallows-type model using the Reverse Major Index (RMJ) as the distance function. They show that, unlike other Mallows-type models, the Mallows-RMJ distribution admits a closed-form expression for the choice probabilities derived from (6):
$\mathbb{P}(y\mid S\,;x)=\tfrac{\phi(x)^{d(y,S)}}{1+\phi(x)+\cdots+\phi(x)^{|S|-1}},$ (9)
where $d(y,S):=\sum_{y^{\prime}\in S\setminus\{y\}}\mathbb{I}\{\mu_{0}^{-1}(y\mid x)>\mu_{0}^{-1}(y^{\prime}\mid x)\}$ equals the number of items in $S$ that are ranked higher than $y$ according to the global ranking $\mu_{0}(\cdot\mid x)$. In other words, $d(y,S)$ defines the relative ranking position of item $y$ within the assortment $S$, so that its choice probability decays exponentially with its rank position in $S$.
@@ 3.3 Gradient Analysis
To gain a deeper mechanistic understanding of the RCPO framework, we analyze the gradient structure of its loss functions. We focus here on the gradient of the Mallows-RMJ-PO-Top-$k$ objective as a representative example. Full derivations and gradient expressions for other cases are deferred to Appendix E. The gradient with respect to model parameters $\theta$ is given by:
@@ 4 Experiments
In this section, we follow a setup similar to that in Meng et al. (2024) and evaluate our methods on widely used benchmarks. Additional experimental details are provided in Appendix F.
@@ 4.1 Experimental Setup
We adopt Llama-3-8B-Instruct and Gemma-2-9B-it as our base models, as both are widely used flagship instruction-tuned models that represent the state-of-the-art. We generate multiple responses to each prompt in the UltraFeedback dataset (Cui et al., 2023) and use the Skywork-Reward-V2-Llama-3.1-8B reward model to provide feedback, which achieves state-of-the-art performance on seven major reward model benchmarks at the time of writing this paper. We refer readers to Appendix F regarding details of constructing the ranking-based preference dataset.
Evaluation. We assess our fine-tuned models on two widely used instruction-following benchmarks: AlpacaEval 2.0 (Dubois et al., 2024) and Arena-Hard-v0.1 (Li et al., 2024). AlpacaEval 2.0 comprises 805 questions drawn from distinct datasets. Performance is measured by the win rate (WR) of model outputs against reference answers generated by GPT-4-Turbo. We also report the length-controlled win rate (LC), which adjusts WR to control for output length. Arena-Hard-v0.1 consists of 500 well-defined technical problem-solving prompts and evaluates models using WR against GPT-4-0314. Previous works (Meng et al., 2024) have demonstrated that Arena-Hard-v0.1 achieves stronger model separability than AlpacaEval 2.0. For both benchmarks, we use GPT-4.1-mini as the judge, replacing the default GPT-4-Turbo due to its improvements (OpenAI, 2025). To enhance cross-judge robustness, we additionally employ GPT-5-mini as the judge on Arena-Hard-v0.1. The results are in Appendix F.1.
4.2 Results

### v2

@@ Abstract
Alignment of large language models (LLMs) has predominantly relied on pairwise preference optimization, where annotators select the better of two responses to a prompt. While simple, this approach overlooks the opportunity to learn from richer forms of human feedback, such as multiway comparisons and top-$k$ rankings. We introduce Ranked Choice Preference Optimization (RCPO), a unified framework that bridges preference optimization with (ranked) choice modeling via maximum likelihood estimation. RCPO supports both utility-based and rank-based models, subsumes several pairwise methods (such as DPO and SimPO) as special cases, and provides principled training objectives for richer feedback formats. We instantiate this framework with two representative models (Multinomial Logit and Mallows-RMJ). Experiments on Llama-3-8B-Instruct, Gemma-2-9B-it, and Mistral-7B-Instruct across in-distribution and out-of-distribution settings show that RCPO consistently outperforms competitive baselines. RCPO shows that directly leveraging ranked preference data, combined with the right choice models, yields more effective alignment. It offers an extensible foundation for incorporating (ranked) choice modeling into LLM training.
@@ 1 Introduction
Large language models (LLMs), a prominent form of generative AI, have rapidly transformed human-computer interaction, powering applications from open-ended dialogue (Thoppilan et al., 2022) and content creation (Brown et al., 2020) to code generation (Chen et al., 2021; Li et al., 2023) and healthcare decision support (Nori et al., 2023). A key driver of this success is alignment: training models to produce outputs that are factual, helpful, safe, and aligned with social norms.
@@
Despite their differences, most alignment methods rely on pairwise preference data, where for each prompt $x$, a preferred response $y_{w}$ and a dispreferred response $y_{l}$ are selected by human annotators or AI judges. In practice, however, preference feedback is often richer than simple pairs. Annotators may provide partial rankings, top-$k$ selections, or single-best judgments from a larger candidate set. Current approaches typically reduce this richer information to pairs—e.g., in training InstructGPT, Ouyang et al. (2022) collected rankings of $K$ responses per prompt but converted them into all $\binom{K}{2}$ pairs; in many academic alignment studies (Meng et al., 2024; Chen et al., 2025; Zhao et al., 2024; Gupta et al., 2025), multiple responses are scored by a reward model, but only the highest- and lowest-scoring are kept. These reductions, while convenient for pairwise-based algorithms, risk distorting the original preference structure and discarding potentially valuable information.
@@
To address this gap, we propose Ranked Choice Preference Optimization (RCPO), a general framework that generalizes pairwise preference alignment to ranked choice feedback (see Figure 1). Rather than restricting evaluators to comparing only two responses, RCPO presents a set of candidates and asks them to choose either the single-best or the top-$k$ responses for a given prompt. Our approach is grounded in the theory of (ranked) choice models, particularly discrete choice modeling, which have been extensively studied in psychology, marketing, economics, and operations research. To the best of our knowledge, RCPO is among the first attempts to systematically apply (ranked) choice modeling to LLM alignment.
@@
(1) Conceptual Framework: We establish a systematic connection between LLM fine-tuning and choice modeling, showing that fine-tuning can be essentially reduced to maximum likelihood estimation (MLE) of choice models. Building on this insight, we develop RCPO as a principled extension of pairwise preference alignment that directly incorporates ranked choice feedback. RCPO is a general and flexible framework: any choice model that satisfies certain regularity conditions can be integrated. By preserving the richness of the original annotations – whether single-best or top-$k$ preferences – RCPO mitigates the information inefficiency in pairwise conversion and enables more effective alignment with human intent.
(2) Concrete Examples: We showcase how two broad classes of choice models, i.e., utility- or rank-based, can be accommodated within the RCPO framework. We then instantiate RCPO with a representative model from each class: the Multinomial Logit (MNL) model (McFadden, 1972) for utility-based choices and the Mallows-RMJ model (Feng and Tang, 2022) for rank-based choices. For both models, we derive alignment objectives under single-best and top-$k$ settings (see Table 1). We also use gradient analysis to shed light on the theoretical underpinnings of these preference optimization methods.
@@
(3) Experiments: We evaluate RCPO on state-of-the-art instruction-tuned LLMs (Llama-3-8B-Instruct, Gemma-2-9B-it, and Mistral-7B-Instruct) using in-distribution test and widely adopted out-of-distribution benchmarks (AlpacaEval 2 and Arena-Hard). Results consistently show that RCPO improves model performance and demonstrates flexibility across base models, preference feedback, and evaluation settings. We highlight the Mallows-RMJ-based preference optimization, which achieves strong results under both pairwise and ranked choice setups.
@@
2 A Choice Based Alignment Framework
@@ 2.1 RLHF and DPO
RLHF comprises three sequential phases. First, it fine-tunes a pre-trained LLM through supervised learning on supervised data and produces an SFT model, which we use as the reference policy $\pi_{\mathrm{ref}}$. Second, RLHF fits a reward model $r^{*}(x,y)$, which can be a neural network itself, based on a separate pairwise preference dataset $\mathcal{D}=\{(x^{(i)},\,y_{w}^{(i)},\,y_{l}^{(i)})\}_{i=1}^{N}$. Here $x^{(i)}$ represents the prompt provided to annotator $i$, and $y_{w}^{(i)}$ and $y_{l}^{(i)}$ are the preferred and dispreferred responses, respectively. Third, the LLM is further fine-tuned via reinforcement learning to maximize the regularized expected reward:
@@
Discrete Choice Models. We also introduce the key concepts for (ranked) choice modeling. First, a rich body of research across economics, marketing, and operations research has developed a variety of discrete choice models to represent human preferences. Let $\mathcal{N}=\{y_{1},y_{2},\ldots,y_{n}\}$ denote the universe of items. A discrete choice model specifies the probability of selecting an item $y$ from an assortment $S\subseteq\mathcal{N}$ under a given context $x$, denoted by $\mathbb{P}(y\mid S\,;x)$. This is an abstraction of many business and economics use cases. For instance, in retail settings, items typically correspond to products, assortments represent the sets of available products at the time of choice, and the context encompasses covariates such as prices, promotions, or product features. There are many ways to define a discrete choice model, which essentially reduce to specifying the probability function $\mathbb{P}(y\mid S\,;x)$.
Ranked Choice Models. Discrete choice models can be extended to richer feedback in the form of ranked choices. We adopt the ranked choice framing of Feng and Tang (2022). Let $\mu^{k}=y_{1}\succ y_{2}\succ\cdots\succ y_{k}$ with $\{y_{1},\ldots,y_{k}\}\subseteq S$ denote a top-$k$ list of items from $S$. A ranked choice model defines $\mathbb{P}(\mu^{k}\mid S\ ;x)$, which specifies the probability of observing the partial ranking $\mu^{k}$ under context $x$. This notion of ranked choices generalizes several common feedback structures, such as (i) pairwise comparisons; (ii) discrete choices (or multiway comparisons); (iii) listwise feedback (or full rankings over the items in any given assortment); (iv) top-$k$ rankings over a full item set, to name a few. For example, when $k=1$, a ranked choice model reduces to a discrete choice model.
@@ 2.3 RCPO: Connection between LLM Alignment and Ranked Choice Modeling
What truly demonstrates the versatility of RCPO is how it gives birth to new preference optimization methods. In this paper, we consider two prominent classes of choice models: utility-based models, which rely on the numerical magnitudes of the items’ utilities, and rank-based models, which depend on the rankings of the items. Despite using different representations of the underlying preference distribution, they are both compatible with the unified RCPO framework.
@@ 2.4 Utility-based Choice Models
RUMs are categorized by the distribution of their stochastic terms. The multinomial logit (MNL) model, introduced by McFadden (1972), assumes i.i.d. Gumbel noise. Alternatives include the probit model (joint normal distribution (Daganzo, 2014)), the nested logit model (correlated extreme value distributions (McFadden, 1980)), and the exponomial model (negative exponential distributions (Alptekinoğlu and Semple, 2016)).
@@ 2.5 Rank-based Choice Models
While utility-based models operate on numerical utilities, rank-based models represent preferences as complete orderings over $\mathcal{N}$, depending only on the relative positions of items [^1] Although technically, an RUM can be transformed into a distribution of rankings, not every RUM can necessarily be transformed into a Mallows-type model. Therefore, we treat them separately in this paper, with different ways to parametrize the choice probabilities.. See Jagabathula and Venkataraman (2022) for a survey of such models.
The Mallows-type model (Mallows, 1957; Fligner and Verducci, 1986) is among the most widely used classes of ranking models. Let $\mathfrak{S}_{n}$ be the set of permutations of $\mathcal{N}$. A Mallows-type model assigns a probability distribution over permutations $\mu\in\mathfrak{S}_{n}$ based on their distance from a central ranking $\mu_{0}$:
@@
where $d(\cdot,\cdot)$ is a distance function between permutations, $\phi\in(0,1)$ is a dispersion parameter. Intuitively, rankings closer to $\mu_{0}$ are exponentially more likely. Different distance choices yield different variants, such as Kendall’s Tau (Mallows, 1957), Spearman’s rank and footrule (Diaconis and Graham, 1977), Hamming distance (Bookstein et al., 2002), Cayley distance (Irurozki et al., 2018), and Reverse Major Index (Feng and Tang, 2022). To capture context dependence, we parameterize the central ranking and dispersion as functions of $x$, denoted $\mu_{0}(\cdot|x)$ and $\phi(x)$. For readability, we may suppress $x$ in the notation when its dependence is unambiguous.
@@ 3.1 Multinomial Logit Model (MNL)
If we take the random shock $\boldsymbol{\varepsilon}$ to be i.i.d. Gumbel, we recover the Multinomial Logit (MNL) model (McFadden, 1972), arguably the most widely used RUM. The corresponding choice probabilities in 4 admit simple closed-form expressions, given by $\mathbb{P}(y_{i}|S;x)=e^{\nu_{y_{i}}(x)}/\sum_{j=1}^{|S|}e^{\nu_{y_{j}}(x)}$. As such, it naturally extends the Bradley-Terry model from pairwise comparisons to multi-item choice settings. By representing the mean utility as the reward defined in 2, we have the following theorem.
@@
As shown in Feng and Tang (2023), the MNL model also extends to a simple ranked choice model. Specifically, the probability that a top-$k$ ranked choice $\mu^{k}=y_{1}\succ y_{2}\succ\ldots\succ y_{k}$ is chosen out of an assortment $S\subseteq\mathcal{N}$ is given by
$\mathbb{P}(\mu^{k}\mid S\ ;x)=\prod\nolimits_{i=1}^{k}\tfrac{e^{\nu_{y_{i}}(x)}}{\sum_{j=i}^{k}e^{\nu_{y_{j}}(x)}+\sum_{y_{h}\in S\setminus\{y_{1},\ldots,y_{k}\}}e^{\nu_{y_{h}}(x)}}.$ (8)
@@ 3.2 Mallows-RMJ Model
A notable challenge in applying Mallows-type models to (ranked) choice modeling is that the ranked choice probabilities in (6) are usually difficult to obtain, since the sum is taken over all permutations. In this regard, an exception is Feng and Tang (2022), who adopt a Mallows-type model using the Reverse Major Index (RMJ) as the distance function. They show that, unlike other Mallows-type models, the Mallows-RMJ distribution admits a closed-form expression for the choice probabilities derived from (6):
$\mathbb{P}(y_{i}\mid S\,;x)=\tfrac{\phi(x)^{d(y_{i},S)}}{1+\phi(x)+\cdots+\phi(x)^{|S|-1}},$ (9)
where $d(y_{i},S):=\sum_{y_{j}\in S\setminus\{y_{i}\}}\mathbb{I}\{\mu_{0}^{-1}(y_{i}\mid x)>\mu_{0}^{-1}(y_{j}\mid x)\}$ equals the number of items in $S$ that are ranked higher than $y_{i}$ according to the global ranking $\mu_{0}(\cdot\mid x)$. In other words, $d(y_{i},S)$ defines the relative ranking position of item $y_{i}$ within the assortment $S$, so that its choice probability decays exponentially with its rank position in $S$.
@@ 3.3 Gradient Analysis
To gain a deeper mechanistic understanding of the RCPO framework, we analyze the gradient structure of its loss functions. We focus here on the gradient of the Mallows-RMJ-PO-Top-k objective as a representative example. Full derivations and gradient expressions for other settings are deferred to Appendix E. The gradient with respect to model parameters $\theta$ is given by:
@@ 4 Experiments
In this section, we assess our methods on several widely used evaluation protocols. Additional experimental details are provided in Appendix F.
@@ 4.1 Experimental Setup
We adopt Llama-3-8B-Instruct, Gemma-2-9B-it and Mistral-7B-Instruct as our fine-tuning bases, as they are widely used flagship instruction-tuned models that represent the state-of-the-art. We generate multiple responses to each prompt in the UltraFeedback train dataset (Cui et al., 2023) and use the Skywork-Reward-V2-Llama-3.1-8B reward model to provide feedback, which achieves state-of-the-art performance on seven major reward model benchmarks at the time of writing this paper (Liu et al., 2025a). We refer readers to Appendix F regarding details of constructing the ranking-based preference dataset.
Evaluation. We evaluate our fine-tuned models on out-of-distribution and in-distribution tests.
For the out-of-distribution test, we assess our models on two widely used instruction-following benchmarks: AlpacaEval 2.0 (Dubois et al., 2024) and Arena-Hard-v0.1 (Li et al., 2024). AlpacaEval 2.0 comprises 805 questions drawn from distinct datasets. Performance is measured by the win rate (WR) of model outputs against reference answers generated by GPT-4-Turbo. We also report the length-controlled win rate (LC), which adjusts WR to control for output length. Arena-Hard-v0.1 consists of 500 well-defined technical problem-solving prompts and evaluates models using WR against GPT-4-0314. Previous works (Meng et al., 2024) have demonstrated that Arena-Hard-v0.1 achieves stronger model separability than AlpacaEval 2.0. For both benchmarks, we use GPT-4.1-mini as the judge, replacing the default GPT-4-Turbo due to its improvements (OpenAI, 2025). To enhance cross-judge robustness, we additionally employ GPT-5-mini as the judge on Arena-Hard-v0.1. The results are in Section F.2.
For the in-distribution test, we first let each fine-tuned model generate responses on the UltraFeedback test dataset, and then compute its win rate against the preferred responses in the dataset, using GPT-4.1-mini as the judge. This methodology aligns with standard evaluation protocols in related literature (e.g., Rafailov et al., 2023; Liu et al., 2023; 2025b).
4.2 Main Results